The mass of the cylinder is 800 g. The heater delivers 8700 J of thermal energy to the cylinder and the temperature of the cylinder increases by 12 °C. Calculate a value for the specific heat capacity of aluminium.


1
Expert's answer
2021-08-30T15:05:38-0400
"Q=mc\Delta T,""c=\dfrac{Q}{m\Delta T},""c=\dfrac{8700\ J}{0.8\ kg\cdot12^{\circ}C}=906\ \dfrac{J}{kg\! \cdot^{\circ}\!C}."